OpenStudy (anonymous):

i need help idk what to do 1.(3a4)3 2.(-2x3y)4 3.(-mn8)2 4.(3x3y)2(-4x2y4)3

OpenStudy (unklerhaukus):

1. \[ (3a^4)^3=(3a^4)(3a^4)(3a^4)=9a^{12}\]

OpenStudy (anonymous):

can u elp with any more?

OpenStudy (unklerhaukus):

2.\[(-2x^3y)^4=(-2x^3y)(-2x^3y)(-2x^3y)(-2x^3y)=\cdots\]

OpenStudy (anonymous):

if there all -2 dose that make them a +?

OpenStudy (unklerhaukus):

\[-2\times-2\times-2\times-2\times=+16\]

OpenStudy (anonymous):

and how would i do 3 and 4 there not the sam as 1 and 2